A faulty thermometer has fixed points at 0.2°c and 101.7°c. What would be the correct reading corresponding to a temperature of 60°c by the thermometer?

Grade:12th pass

3 Answers

Arun
4 years ago
 
Here faulty reading (x)=60
Lower fixed point of faulty readind (L.F.P)=0.2
Upper fixed point of faulty reading (U.F.P.)=101.7
Fundamental interval (F.I.)=101.7 – 0.2 = 101.5
Correct reading(C)=?
c/100*(x-L.F.P)/F.I.
C/100=59.8/101.5 = 0.589
C=0.589*100=58.9 celsius scale
Khimraj
4 years ago
usually the thermometer fixed points basically are:
Lower point =0°c
Upper point=100°c
so in between there are 100 divisions.
each one division corresponds to 1°c
But in the above mentioned case :
Lower point=0.2°c
upper point=101.7°c
so total division=101.7-0.2 = 101.5
thus, 1 degree= 1 division=100/101.5
Now let us find the given temperature:
measured temperature 60 °c
=100x60/101.5
=59.11°c
Hope it is clear.
Rajdeep
4 years ago
HELLO THERE!
 
Here’s the generalized formula:
 
Let C be the correct reading,
be the lower point of faulty thermometer,
be the upper point of faulty thermometer,
be the given faulty reading.
 
\frac{C}{100} = \frac{x - L}{U - L} \\\\\implies \frac{C}{100} = \frac{60 - 0.2}{101.7 - 0.2} \\\\\implies C = 58.9^{o}C
 
THANKS!